# A Stochastic Approach to the Definition of the Path Integral Measure

### v1

@@ 2.2 Hamilton’s Principle and Boundedness
Proof: Let $X$ be the dual to $\mathcal{P}$. Given $S:\mathcal{P}\rightarrow\mathbb{R}$ with $S\in C^{2}$ and weakly-differentiable, $\exists\mathcal{A}\subset E:\{S^{\alpha}_{\gamma}\}$ defining the set of all action functionals for transformations of $\gamma_{E}$ defining elements of $\mathcal{A}$. Then, with $S_{1}\preceq S_{2}:S_{1}:=S(\gamma_{E}^{\alpha_{1}}),S_{2}:=S(\gamma_{E}^{\alpha_{2}}),\alpha_{1}\leq\alpha_{2}$. It follows that $|S_{2}(\gamma)-S_{1}(\gamma)|\leq|S_{2}(\gamma)-S_{2}(\gamma_{E})|\quad\forall\gamma\in\mathcal{T}_{r}$, given that $S_{2}(\gamma)=S_{1}(T_{\alpha_{2}}T_{\alpha_{1}}\gamma^{\alpha_{0}}_{0})=T_{\alpha_{2}}S_{1}(\gamma)$. We obtain a chain $\{S_{\alpha}\}\subset\mathcal{A}$ with a pointwise upper bound. This limit has the same Euler-Lagrange equation as all of the actions, and thus by Zorn’s lemma, $\exists\bar{S}\in\mathcal{A}$ with $\eta=\sup_{\gamma\in\mathcal{T}_{r}}|\bar{S}(\gamma)-\bar{S}(\gamma_{E})|.$ From maximality, no other admissible action can be greater in difference than $\eta$, and therefore, $\exists\mathcal{T}_{\eta}\in\mathcal{P}:\forall\gamma$ with energy in $E-\delta E,E+\delta E]$, $\gamma\in\mathcal{T}_{\eta}.$
@@ 3.1 Geometry of the $\eta$-Tube and Bundle structure
The flux space along $\gamma$ is the Bochner $L^{2}$–space $F$ defined in (3.3), with $u(t)\in T_{\gamma(t)}M$ and $\lvert u(t)\rvert_{g}^{2}:=g_{\gamma(t)}\bigl(u(t),u(t)\bigr)$. The inner product on $F$ is $\langle u,v\rangle_{L^{2}}:=\int_{0}^{1}g_{\gamma(t)}\bigl(u(t),v(t)\bigr)dt.$ Elements of $F$ are fluxes, i.e. square–integrable vector fields along $\gamma$.
$F:=L^{2}\bigl([0,1];\gamma^{*}TM\bigr):=\Bigl\{u:[0,1]\to\gamma^{*}TM\ \text{measurable}:\ \int_{0}^{1}\lvert u(t)\rvert_{g}^{2}\,dt<\infty\Bigr\},$ (3.3)
@@ 4.1 Formulating the Measure and the Stochastic Path Integral
On $TM$ with $\pi:E\rightarrow[0,1],\quad E=\cup_{t\in[0,1]}T_{\gamma(t)}M,\quad\exists\mu^{TM}:\forall A$—measurable, $\mu^{TM}=\int_{0}^{1}\mu_{t}(A)\bar{\rho}(dt),$ where $\bar{\rho}=\pi_{*}\mu^{TM}$ a conditional fiber and $\quad\mu^{TM}$—a $\sigma$- additive probability on $E$ and $A$ is Borel, belonging to the Borel $\sigma$-algebra $\mathcal{B}([0,1],M)$. For each $t\in[0,1]$, let $\mu_{t}:=(\mathrm{ev}_{t})_{*}\mu,\qquad\mu_{t}(A)=\mu\big(\{\omega\in\mathcal{C}:\ X_{t}(\omega)\in A\}\big),$ which define the tt—marginal on the fibre $T_{\gamma(t)}M$. Let $E:=\bigsqcup_{t\in[0,1]}T_{\gamma(t)}M$ be the bundle with projection $\pi:E\to[0,1]$, and let $\bar{\rho}$ be a probability measure on $[0,1]$, which is the bundle we have defined for the flux tubes.
@@
This theorem will not be proven, and may even be understood better as a definition rather than the former. However, we will maintain it as being a theorem given that we will reference it constantly in the proofs of equivalence with the Euclidean Path Integral rigorously defined in the Feynman-Kac formalism. The rest of the current section will be dedicated to establishing the properties of the Stochastic Path integral. We would also like to underline that the phase factor $\exp(\frac{i}{\hbar}S)$ is posited. While it is possible to derive it from principles of statistical mechanics in application to our setup, this will not be performed here.

### v2

@@ 2.2 Hamilton’s Principle and Boundedness
[…]
@@ 2.3 The Quantum-Mechanical View
3 On Relevant Trajectories
In order to define what trajectories we can account for under an uncertainty in energy, we ought to define a mathematical criterion that could quantitatively assess whether a trajectory is allowable–as in the energy uncertainty permits a classical particle to deviate from the classical trajectory by this alternative–or whether that trajectory is forbidden.
Let us denote the energy uncertainty by $\delta E$ and our probe function by $\Phi$. Given that it is irrelevant whether this function is defined with any specific orientation, we could select it as being defined on a contour formed the classical trajectory $\gamma_{0}$ and any $\gamma$ in its homotopy class. Naturally, under the supposition that the endpoints of the trajectories are unaltered, the two are sufficient to form a closed contour on path space. We will define $\Phi(\gamma,\gamma_{0}):\forall|\Delta E|>\delta E,\Phi\rightarrow\infty$, while for $|\Delta E|\leq\delta E$, $0<\Phi<\infty$.
A Natural choice for such a function would be a contour integral with the integrand that is a meromorphic function that has poles at $E_{0}\pm\delta E.$ For such a function, we can write
$f(E)=\frac{1}{(\delta E)^{2}-(\Delta E)^{2}}\implies f(H)=\frac{1}{(\delta E)^{2}-(H-E_{0})^{2}}$ (3.1)
The corresponding probe function on can be chosen to also depend on the canonical 1-form to be translation-invariant, thus having $\Phi$ vanish for paths that have a purely gauge nature. We obtain
$\Phi(\gamma,\gamma_{0})=\oint_{\gamma\circ\gamma_{0}}f(H)p_{i}dq^{i}$ (3.2)
For any $\gamma:|H(\gamma)-E_{0}|>\delta E,$ the contour approaches a pole of $f(H)$, wherefrom by the Residue Theorem $\Phi(\gamma,\gamma_{0})$ diverges. The class of admissible trajectories is therefore given by the homotopies of $\gamma_{0}$ for which $\Phi$ is finite.
@@ 4.1 Geometry of the $\eta$-Tube and Bundle structure
The flux space along $\gamma$ is the Bochner $L^{2}$–space $F$ defined in (3.3), with $u(t)\in T_{\gamma(t)}M$ and $\lvert u(t)\rvert_{g}^{2}:=g_{\gamma(t)}\bigl(u(t),u(t)\bigr)$. The inner product on $F$ is $\langle u,v\rangle_{L^{2}}:=\int_{0}^{1}g_{\gamma(t)}\bigl(u(t),v(t)\bigr)dt.$ Elements of $F$ are fluxes, i.e. square–integrable vector fields along $\gamma$ for measurable $u:[0,1]\rightarrow\gamma^{*}TM$ and
$F:=L^{2}\bigl([0,1];\gamma^{*}TM\bigr)=\Bigl\{u:\ \int_{0}^{1}\lvert u(t)\rvert_{g}^{2}\,dt<\infty\Bigr\},$ (4.3)
@@ 4.2 Stochastic Dynamics in $\mathcal{T}_{\eta}$
It is important to discuss in more detail the trajectory fluctuations that are here being accounted for. If we choose to consider the fluctuations belonging to the normal frame $N_{\gamma_{0}}M$ over any $V^{n}$, the dimensionality of the contribution of the fluctuations to the final propagator will be equal to n1n-1. This can be seen by computing (4.5) for a free particle, which yields $(\frac{m}{2\pi i})^{\frac{n-1}{2}}$ rather than the standard n-dimensional result from the traditional path integral. Luckily, this issue can be solved by taking into account longitudinal fluctuations.
So far, we have been considering the fluctuations $X_{t}$ as being defined over the normal bundle, which would yield an orthogonal decomposition roughly of the form $X(t)=\gamma_{0}(t)+p_{N}(t):p_{N}\in N_{\gamma_{0}}$. Now, if we consider longitudinal fluctuations $p_{L}(t):p_{L}(0)=p_{L}(1)=0,$ we could correct our fluctuation to be $X_{t}=\gamma_{0}(t)+p_{N}(t)+p_{L}(t)\dot{\gamma}_{0}(t)$. Under infinitesimal reparameterization on the Brownian bridge $t\rightarrow t+\epsilon(t):\epsilon(0)=\epsilon(1)=0,$, we obtain that $\delta X(t)=\epsilon(t)\dot{\gamma}(t)\forall\gamma\in\text{Hom}\gamma_{0}$, wherefrom longitudinal variations are gauges. Therefore, we find ourselves in need of removing the zero mode. This may be done by introducing a gauge condition
$G(p_{L})=\int_{0}^{1}p_{L}(t)dt=0,$ (4.13)
We may then introduce the FP identity, by which we get (3.14), removing the zero mode.
$\int\delta(G(p_{L}+\epsilon))\det(\frac{\delta G}{\delta\epsilon})\mathcal{D}\epsilon=1$ (4.14)
@@ 5.1 Formulating the Measure and the Stochastic Path Integral
On $TM$ with $\pi:E\rightarrow[0,1],\quad E=\cup_{t\in[0,1]}T_{\gamma(t)}M,\quad\exists\mu^{TM}:\forall A$—measurable, $\mu^{TM}=\int_{0}^{1}\mu_{t}(A)\bar{\rho}(dt),$ where $\bar{\rho}=\pi_{*}\mu^{TM}$ a conditional fiber and $\quad\mu^{TM}$—a $\sigma$- additive probability on $E$ and $A$ is Borel, belonging to the Borel $\sigma$-algebra $\mathcal{B}([0,1],M)$. For each $t\in[0,1]$, let $\mu_{t}:=(\mathrm{ev}_{t})_{*}\mu,\qquad\mu_{t}(A)=\mu\big(\{\omega\in\mathcal{C}:\ X_{t}(\omega)\in A\}\big),$ which define the tt—marginal on the fibre $T_{\gamma(t)}M$. Let $E:=\bigsqcup_{t\in[0,1]}T_{\gamma(t)}M$ be the bundle with projection $\pi:E\to[0,1]$, and let $\bar{\rho}$ be a probability measure on $[0,1]$, which is the bundle we have defined for the flux tubes. The Gaussian measure $\Gamma$ splits into a longitudal and normal measures, such that $\Gamma_{N}\in H^{1}_{0}([0,1],\mathbb{R}^{n-1})$ and $\Gamma_{L}\in H^{1}_{0}([0,1])$, with $\Gamma=\Gamma_{N}\otimes\Gamma_{L}$.
@@
The rest of the current section will be dedicated to establishing the properties of the Stochastic Path integral. We would also like to underline that the phase factor $\exp(\frac{i}{\hbar}S)$ is posited. While it is possible to derive it from principles of statistical mechanics in application to our setup, this will not be performed here.